# post-last exit time process and its application to Loss-Given-Default Distribution

### v3

@@ 2.2. Post-last exit time process
For the post-exit time process, we need only to consider a point $y<\alpha$. In view of (2.14), (2.5) becomes
This means that the post-last exit time process is obtained equivalently via the transform $g(\cdot):=s(\alpha)-s(\cdot)$. Call this transform also $X^{h_{\alpha}}$ and its scale function and speed measure $s^{g}$ and $m^{g}$, respectively. By the general result for the transform of this kind (see Appendix A.1), we have
$=\int_{y}^{\alpha}(s(\alpha)-s(\eta))m({\rm d}\eta)<\infty.$
The last finiteness result is due to $\alpha$ being a regular point for $X$. This proves that $\alpha$ is an entrance boundary for the $g$-transform, which is equivalent to the post-exit time process $X^{h_{\alpha}}$ (see [Karlin and Taylor (1981), Chap.15, Table 6.2]). ∎
2.3. Excursion-theoretic view
We can view $\{X_{t},t>L_{\alpha}\}$ as an excursion from $\alpha$ of infinite length. To pursue this point of view, we review some of the excursion theory based on Getoor (1979) and Rogers and Williams (1994b), in particular its Sections VI. 50, 54 and 55.
Let us return to the original process $X$ and consider an excursion from a regular point $\alpha$. Let the local time at $y\in\mathcal{I}$ be $l_{t}^{y},t\geq 0$. Set $l_{t}=l_{t}^{\alpha}$ and define the right inverse $\rho(t,\alpha):=\inf\{s:l_{s}>t\}$. It is well known that $\rho$ is a right-continuous and nondecreasing subordinator under $\mathbb{P}_{\alpha}$ and hence
$\mathbb{E}_{\alpha}\left[e^{-q\rho(t,\alpha)}\right]=e^{-tg(q)}$ (2.16)
where
$g(q)=m\{\alpha\}q+\int_{(0,\infty]}(1-e^{-qs})\nu({\rm d}s)$ (2.17)
By using the tail distribution
$c(u):=\nu((u,\infty]),$
the Laplace exponent is rewritten as
$g(q)=m\{\alpha\}q+q\int_{0}^{\infty}e^{-qs}c(s){\rm d}s.$ (2.18)
We define
$r_{q}(x,y):=\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qt}{\rm d}l^{y}_{t}\right),\quad x,y\in\mathcal{I},q>0.$
Then, by a change of variable and (2.16),
$r_{q}(\alpha,\alpha)=\mathbb{E}_{\alpha}\left(\int_{0}^{\infty}e^{-qt}{\rm d}l_{t}\right)=\mathbb{E}_{\alpha}\left(\int_{0}^{\infty}e^{-q\rho(t,\alpha)}{\rm d}t\right)=\frac{1}{g(q)}.$ (2.19)
$U^{q}f(x)=\int_{\mathcal{I}}m({\rm d}y)f(y)\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qs}{\rm d}l^{y}_{s}\right)$
which is also written as
$U^{q}f(x)=\int_{\mathcal{I}}m({\rm d}y)f(y)G_{q}(x,y)$
by using $G_{q}(x,y)=\int_{0}^{\infty}e^{-qt}p(t;x,y){\rm d}t$. By comparison, $G_{q}(x,y)=\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qs}{\rm d}l^{y}_{s}\right)$ a.e. $y$. But by the right continuity in $y$, we have the equality for every $y\in\mathcal{I}$. We obtain from (2.19) that $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$ and hence
Remark 2.1.
Let $X$ be a linear diffusion with regular point $\alpha\in\mathcal{I}$. We consider the local time at $\alpha$ and excursions from this point. With $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$, (2.18) becomes
$\int_{0}^{\infty}e^{-qs}c(s){\rm d}s=\frac{1}{qG_{q}(\alpha,\alpha)},$
which can be used to obtain $c(u)=\nu((u,\infty])$ through the inverse Laplace transform. This $c(u)$ corresponds to the rate at which the first excursion of length greater than $u$ occurs. To consider an excursion of infinite length, which is closely related to a post-last exit time process discussed in Section 2.2, let us take $q\downarrow 0$ in (2.16) to find
$\mathbb{P}_{\alpha}(\rho(t,\alpha)<\infty)=e^{-t\nu(\{+\infty\})}$
by (2.17). Moreover, from $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$, we have
$\nu(\{+\infty\})=\lim_{q\downarrow 0}\frac{1}{G_{q}(\alpha,\alpha)}:=\frac{1}{G_{0}(\alpha,\alpha)},$ (2.20)
from which we define
$A(x):=\mathbb{P}_{\cdot}(l_{L_{\alpha}}\leq x)=1-e^{-\frac{x}{G_{0}(\alpha,\alpha)}}$
Hence with $95\%$ probability, $l_{L_{\alpha}}$ occurs before $A^{-1}(0.95)$. $\diamondsuit$
Note that the quantity $G_{0}(y,z)$ for $y,z\in\mathcal{I}$ has a representation in terms of the scale function:
$G_{0}(y,z)=\int_{0}^{\infty}p(t;y,z){\rm d}t=\begin{cases}\lim\limits_{a\downarrow\ell,b\uparrow r}\frac{(s(y)-s(a))(s(b)-s(z))}{s(b)-s(a)},\quad\ell<y\leq z<r,\\ \lim\limits_{a\downarrow\ell,b\uparrow r}\frac{(s(z)-s(a))(s(b)-s(y))}{s(b)-s(a)},\quad\ell<z\leq y<r.\end{cases}$ (2.21)
Another important element of the excursion is the entrance law of the excursion measure,
$\mathfrak{n}^{\alpha}_{t}(\cdot)=\mathfrak{n}^{\alpha}(\{\varepsilon:\varepsilon(t)\in\cdot,t<\zeta^{\prime}(\epsilon)\}),$
where $\mathfrak{n}^{\alpha}$ is the characteristic (or Itô) measure of the excursion process, $\epsilon(t)$ is a generic excursion process from point $\alpha$, and $\zeta^{\prime}$ is lifetime of the excursion. The following is a simple but, to the best of our knowledge, a general result since it does not resort to the symmetry of transition densities (see Csáki et al. (1987)):
Proposition 2.3.
Let $X$ be a linear diffusion with regular points $\alpha,y\in\mathcal{I}$ and $m({\rm d}y)=m(y){\rm d}y$ be its speed measure. Assume $\alpha>y$. Then, the entrance law $\mathfrak{n}^{\alpha}_{t}({\rm d}y)=\mathfrak{n}^{\alpha}_{t}(y){\rm d}y$ is given by
$\mathfrak{n}^{\alpha}_{t}(y)=q(t;\alpha,y)m(y),\quad\text{a.e.}\>y\in(\ell,\alpha)$ (2.22)
where $q(t;\alpha,y)$ is the density of the first passage time distribution from $y$ to $\alpha$: $\mathbb{P}_{y}(H_{\alpha}\in{\rm d}t)=q(t;\alpha,y){\rm d}t$.
Proof.
The entrance law has the characterization (see e.g. VI.(50.3) in Rogers and Williams (1994b))
$\int_{0}^{\infty}e^{-qt}\mathfrak{n}^{\alpha}_{t}(\Gamma){\rm d}t=\frac{U^{q}\mbox{1}\mbox{l}_{\Gamma}(\alpha)}{\mathbb{E}_{\alpha}\left(\int_{0}^{\infty}e^{-qt}{\rm d}l_{t}\right)}$
where $\Gamma$ is a Borel set not containing $\{\alpha\}$. By (2.4) and (2.19),
$\int_{0}^{\infty}e^{-qt}\int_{\Gamma}\mathfrak{n}^{\alpha}_{t}(y){\rm d}y{\rm d}t$ $=\int_{0}^{\infty}e^{-qt}\mathfrak{n}^{\alpha}_{t}(\Gamma){\rm d}t=\frac{\int_{\Gamma}G_{q}(\alpha,y)m({\rm d}y)}{G_{q}(\alpha,\alpha)}=\int_{\Gamma}\frac{\phi_{q}(\alpha)\psi_{q}(y)}{\phi_{q}(\alpha)\psi_{q}(\alpha)}m({\rm d}y).$
Hence by Fubini, we have for almost every $y$ in $(\ell,\alpha)$,
$\int_{0}^{\infty}e^{-qt}\mathfrak{n}^{\alpha}_{t}(y){\rm d}t=\frac{\psi_{q}(y)}{\psi_{q}(\alpha)}m(y).$ (2.23)
Now by (2.3) in the second equality below,
$\int_{0}^{\infty}e^{-qt}\mathfrak{n}^{\alpha}_{t}(y){\rm d}t=\frac{\psi_{q}(y)}{\psi_{q}(\alpha)}m(y)=\mathbb{E}_{y}[e^{-qH_{\alpha}}]m(y)=\int_{0}^{\infty}e^{-qt}\mathbb{P}_{y}(H_{\alpha}\in{\rm d}t)m(y),$
from which we obtain (2.22). ∎
Remark 2.2.
Note that for the computation of the entrance law, the inverse Laplace transform based on (2.23) also works. $\diamondsuit$
The next object is the entrance law associated with an excursion of infinite length: $\mathfrak{n}^{\alpha}(\{\varepsilon:\varepsilon(t)\in\cdot,|\zeta^{\prime}(\epsilon)=\infty\})$.
Corollary 2.1.
The entrance law of excursions below $\alpha$ of the post-last exit time process $(X_{t})_{t>L_{\alpha}}$ is given by
$v_{t}^{\alpha}(y)=G_{0}(\alpha,\alpha)\cdot h_{\alpha}(y)q(t;\alpha,y)m(y){\rm d}y,\quad y<\alpha.$
Proof.
By setting $E_{\infty}:=\{\epsilon:\zeta^{\prime}(\epsilon)=\infty\}$, it follows that
$\mathfrak{n}^{\alpha}(\{\varepsilon:\varepsilon(t)\in{\rm d}y,|\zeta^{\prime}(\epsilon)=\infty\})$ $=\frac{\mathfrak{n}^{\alpha}(\{\varepsilon:\varepsilon(t)\in{\rm d}y,\zeta^{\prime}(\epsilon)=\infty\})}{\mathfrak{n}^{\alpha}({E_{\infty}})}$
$=\frac{\mathfrak{n}^{\alpha}(\zeta^{\prime}(\epsilon)=\infty|\epsilon(t)\in{\rm d}y)}{\mathfrak{n}^{\alpha}({E_{\infty}})}\mathfrak{n}^{\alpha}(\epsilon(t)\in{\rm d}y)$
$=\frac{\mathbb{P}_{y}(L_{\alpha}=0)}{\nu(\{+\infty\})}\mathfrak{n}^{\alpha}_{t}({\rm d}y)$
$=G_{0}(\alpha,\alpha)\cdot h_{\alpha}(y)q(t;\alpha,y)m(y){\rm d}y$
by (2.20) and (2.22). ∎
Figure 1. A typical path of the original process $X$
2.4. Reversed process
Assumption 1 is still in place. Figure 1 shows a typical path of the original process $X$. We wish to identify the reversed process of the $h_{\alpha}$-transform. Recall that $r$ is the right-end point of the state space. A minimal excessive function $k_{r}(x):=s(x)-s(\ell)$ produces the $k_{r}$-transform $X^{r}$: $\mathbb{P}^{k_{r}}_{x}(\lim_{t\rightarrow\zeta}\omega(t)=r)=1$.
Proposition 2.4.
Under Assumption 1 and $H_{\ell}<\infty$ almost surely, the reversed process of the $h_{\alpha}$-transform has the same distribution as the process starting at $\ell$, conditioned to hit $r$, and killed at the first hitting time to $\alpha$:
[…]
are identical in law.
Proof.
Let us first concentrate on the entire path: from time 00 via $L_{\alpha}$ to $H_{\ell}$. It is shown (see [Egami and Kevkhishvili (2020), Proposition 3.6]) that
$A:=\{\omega(H_{\ell}-t),0<t<H_{\ell},\mathbb{P}_{x}\}\quad\text{and}\quad B:=\{\omega(t),0<t<L_{x},\mathbb{P}^{k_{r}}_{\ell}\}$ (2.24)
are identical in law and $\ell$ is an entrance boundary for $\mathbb{P}^{k_{r}}_{\ell}$.
Next we concentrate on the dashed part of the path in Figure 1: from time 00 to $L_{\alpha}$. Recall that $k_{\alpha}(y):=\mathbb{P}_{y}(L_{\alpha}>0)$ is another minimal excessive function and the original process $X$ transformed by this function $k_{\alpha}$ (called the $k_{\alpha}$-transform) is identical in law with $X$ conditioned to hit $\alpha$ and killed at the last exit time from $\alpha$. It is known (see [Salminen (1984), Remark 9 (ii)]) that
$a:=\{\omega(L_{\alpha}-t),0<t<L_{\alpha},\mathbb{P}_{x}\}\quad\text{and}\quad b:=\{\omega(t),0<t<L_{x},\mathbb{P}^{k_{r}}_{\alpha}\}$ (2.25)
are identical in law. Hence the reversed process of the $k_{\alpha}$-transform has the same distribution as the process starting at $\alpha$, conditioned to hit $r$, and killed at the last exit time from $x$.
Now we turn to the red part of the path in Figure 1: from time $L_{\alpha}$ to $H_{\ell}$. Namely, the path of the post-last exit time process $X^{h_{\alpha}}_{t}(\omega)=\{\omega(t):L_{\alpha}<t<H_{\ell},\mathbb{P}^{h_{\alpha}}_{\alpha}\}$. Let us denote the processes in the statement of this proposition by
$a^{\prime}:=\{\omega(H_{\ell}-t),L_{\alpha}<t<H_{\ell},\mathbb{P}_{\alpha}\}\quad\text{and}\quad b^{\prime}:=\{\omega(t),0<t<H_{\alpha},\mathbb{P}^{k_{r}}_{\ell}\}.$
By Corollary 11.30 of Chung and Walsh (2004), the laws of $A$ and $a$ are the same. Since a sample path of $a^{\prime}$ is part of a sample path of $A$ (we are talking about the original path, not the transforms), the laws of $a$ and $a^{\prime}$ are the same. Moreover, by (2.24) and (2.25), so are the laws of $B$ and $b$ and hence of $b$ and $b^{\prime}$. This implies the equivalence of the laws of $a^{\prime}$ and $b^{\prime}$. ∎
Remark 2.3.
(i) Note that unlike (2.24) or (2.25), neither the law of $a^{\prime}$ nor $b^{\prime}$ involves time reversal from co-optional time. Therefore, the celebrated Nagasawa’s theorem employed in Sharpe (1980) cannot be used here for the purpose of identifying a reversed process: a proof as in Proposition 2.4 is necessary.
(ii) We have seen that $\alpha$ is an entrance boundary for the $h_{\alpha}$-transform (Proposition 2.2) and $\ell$ is for the $k_{r}$-transform that starts at $\ell$. This is well reflected in the fact that
$s^{h_{\alpha}}(x)=s^{g}(x)=\frac{1}{s(\alpha)-s(x)}\quad\text{and}\quad s^{r}(x)=-\frac{1}{s(x)-s(\ell)},$
where we notice that $\alpha$ and $\ell$ are interchanged. It is a natural result in view of Proposition 2.4$\diamondsuit$
We summarize this section in Table 1:
Table 1. Let $X$ be a transient linear diffusion with $s(\ell)>-\infty,s(r)=+\infty$. For a regular point $\alpha$, set $h_{\alpha}(x):=\mathbb{P}_{x}(L_{\alpha}=0)$. The entries are the transition density with respect to $m(\cdot)$, the entrance law of excursions below $\alpha$, and the diffusion having the same distribution as the respective reversed process.
Transition density Entrance law Reversed process
Original process $p(t;x,y)$ $n^{\alpha}_{t}(y)=q(t;\alpha,y)m(y)$$\{\omega(t),0<t<L_{x},\mathbb{P}^{k_{r}}_{\ell}\}$
Post-last exit time process $\frac{h_{\alpha}(y)}{h_{\alpha}(x)}p(t;x,y)$ $G_{0}(\alpha,\alpha)h_{\alpha}(y)n^{\alpha}_{t}(y)$$\{\omega(t),0<t<H_{\alpha},\mathbb{P}^{k_{r}}_{\ell}\}$
3. Application to loss-given-default distribution
We propose, as a financial application of post-last exit time process, a model to estimate loss-given-default (LGD) distribution of corporate debt. More specifically, we derive the distribution of the ratio of the firm’s total assets over its total debt upon default. We shall hereafter refer to this ratio as the leverage ratio and to the dynamics of this ratio as the leverage ratio process, denoted by $Y$.  [^1] In our framework, we are focusing on the loss rate of the debt of the whole company, not each individual debt obligation. Based on this corporate-level loss distribution, credit managers can calculate loss rates for individual debt obligations having distinct characteristics; for example, senior or subordinated, and with or without collateral.
In the firm-value (or Merton (1974)) model of credit risk, insolvency occurs upon the leverage ratio being equal to some level $c$. In this case, if we set $\alpha>c$, then there must exist the last exit time $L_{\alpha}$ recorded by the leverage ratio process. The idea is summarized as follows:
  1. (1)
  2. (2)
    Since $L_{\alpha}$ is not a stopping time, it is appropriate to model the interval, denoted by $\tau$, between $L_{\alpha}$ and default time $\xi$ by the intensity-based model:
    $\xi(\omega):=L_{\alpha}(\omega)+\tau\circ\theta(L_{\alpha}(\omega)).$
    In fact, this well explains the reality that a default occurs surprisingly during the period when non-insider investors do not have enough information about credit quality. The $L_{\alpha}$, not a stopping time, corresponds to the time when creditworthiness becomes hard to grasp, and $\xi$ occurs as a surprise after non-predicting $\tau$ has elapsed.
  3. (3)
    The company’s leverage ratio at $\xi$ is then computed as the distribution of $Y(\tau\circ\theta(L_{\alpha}))$, from which the LGD distribution at the corporate level is obtained. For this purpose, we use results from Section 2 including (2.13).
We explicitly derive the distribution of default time $\xi$ (Remark 3.1) and LGD (Corollary 3.2). These are computationally simple formulas. Furthermore, the model check conducted in Section 4 demonstrates that the model-implied default time and LGD distribution are consistent with the credit market and company-specific financial conditions. Specifically, the CDS spread implied from the estimated LGD distribution is consistent with the quoted spread in the market.
3.1. Leverage ratio process
We define the leverage ratio process and compute its infinitesimal drift and diffusion parameters by Lemma 2.1. Let us assume that the firm value (market value of total assets) $V$ follows a geometric Brownian motion:
${\rm d}V_{t}=\mu V_{t}{\rm d}t+\sigma V_{t}{\rm d}W_{t}$ (3.1)
with $\mu\in\mathbb{R}$ and $\sigma>0$. The debt process is given by
$B_{t}=B_{0}e^{rt},\quad B_{0}\in\mathcal{F}_{0},\quad t\geq 0,$
where $r>0$ denotes a constant rate. As announced above, we define the leverage ratio process $Y=V/B$:
The parameters $\mu$, $\sigma$ are such that the left boundary 00 is attracting and the right boundary +$\infty$+ ∞ is non-attracting, which comply with Assumption 1. In particular, we assume that $\mu-\frac{1}{2}\sigma^{2}-r<0$. This ensures that the default occurs eventually with probability one. By fixing a certain level $\alpha\in\mathbb{R}_{+}$, we define
$L_{\alpha}:=\sup\{t\geq 0:Y_{t}=\alpha\}.$
We have the normalized process
by setting
$M:=\frac{\mu-\frac{1}{2}\sigma^{2}-r}{\sigma}.$
Now let us introduce the post-last exit time process
$Z_{t}:=\left(\frac{1}{\sigma}\ln(Y_{t})\right)\circ\theta(L_{\alpha}),\quad t>0.$
Using (2.15) and the scale function $s(y)=\frac{1}{2M}(1-e^{-2M\cdot y})$ of the Brownian motion with drift $M$, we see that the process $Z_{t}$ follows the dynamics
@@ 3.1. Leverage ratio process
The starting point of $Z$ is $\alpha^{*}$. Note that since $\lim_{x\uparrow 0}\coth(x)=-\infty$, when the process $Z$ approaches $\alpha^{*}$ from below, the drift approaches negative infinity so that the process shall never reach the point $\alpha^{*}$ from the region $(-\infty,\alpha^{*})$. Mathematically, as proved in Proposition 2.2, $\alpha^{*}$ is an entrance boundary for $Z$ and financially, the firm’s leverage ratio process $Y$ shall never recover to the threshold level $\alpha$.
@@ 3.2. Loss-given-default distribution
We define a random time that is suitable for default time. As is seen in the intensity-based default modeling, we need an exponential random variable $J$ with rate $\eta=1$, independent of everything else. We let $\lambda:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ be a nonnegative piecewise continuous function and define the moment inverse of integral functional by
$\nu(J):=\min\left(s:\int_{0}^{s}\lambda\left(Z_{u}\right){\rm d}u=J\right).$ (3.4)
@@
We assume that $\lambda$ satisfies
$\int_{0}^{\infty}\lambda\left(Z_{s}\right){\rm d}s=\infty\quad\mathbb{P}-a.s.$ (3.5)
$\diamondsuit$
@@
$\xi(\omega):=L_{\alpha}(\omega)+\tau\circ\theta(L_{\alpha}(\omega))$ (3.6)
where $\theta(\cdot)$ denotes the shift operator. Note that the construction of the random time is similar to the canonical construction of the default time in the intensity-based approach (see Proposition5.26 in Capiński and Zastawniak (2017)). It follows from (3.4) that $\mathbb{P}\left(\tau=0\mid Z_{0}=\alpha^{*}\right)=0$ and from (3.5) that $\mathbb{P}\left(\tau<+\infty\mid Z_{0}=\alpha^{*}\right)=1$. We also see that $\mathbb{P}\left(\tau>s\mid Z_{0}=\alpha^{*}\right)>0$ for every $s\geq 0$. Moreover, we have $\mathbb{P}(L_{\alpha}<\xi<+\infty)=1$ with (3.6). We have constructed the time $\tau$ using the process which never returns to $\alpha^{*}$, this feature being represented in (3.3).
@@
Due to $Y_{\xi}=Y_{\tau}\circ\theta(L_{\alpha})$, it suffices to compute the distribution of $Z_{\tau}=\frac{1}{\sigma}\ln(Y_{\tau}\circ\theta(L_{\alpha}))$. Thus, we are able to find the LGD distribution by focusing on $Z_{\tau}$ only. This simplifies the analysis as we do not need to consider the shift operator $\theta$ in the definition (3.6).
Our setup is summarized in Table 2. For the LGD modeling, we have made only Assumptions 1 and 2.
@@
[…]
@@
$U(z):=\mathbb{E}_{z}\left[e^{-\gamma\tau}\mbox{1}\mbox{l}_{\{Z_{\tau}\leq Q\}}\right],\quad z<\alpha^{*},\quad\gamma>0$ (3.8)
@@
$\frac{1}{2}U^{\prime\prime}(z)+M\coth(M(z-\alpha^{*}))U^{\prime}(z)-(\lambda(z)+\gamma)U(z)=-\lambda(z)\mbox{1}\mbox{l}_{[-\infty,Q]}(z),\quad z<\alpha^{*}.$ (3.9)
Here $\mathbb{E}_{z}[\cdot]$ denotes the conditional expectation $\mathbb{E}[\cdot\mid Z_{0}=z]$. We will provide an explicit form of $U(z)$ in Proposition 3.1. Using the solution $U(z)$, we obtain $\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]=\lim_{z\uparrow\alpha^{*}}U(z)$ by the continuity of $U(z)$. Finally, $\lim_{\gamma\downarrow 0}\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]$ produces the distribution of $K^{B}(\xi)$ via (3.7).
To obtain a concrete result, we specify the function $\lambda$ in (3.4) as
$\lambda\left(Z_{t}\right):=\mbox{1}\mbox{l}_{\{Z_{t}<\alpha^{*}\}}$ (3.10)
so that the integral in (3.4) represents the occupation time of the process $Z$ under the level $\alpha^{*}$.
Hence we identify the firm’s default once the leverage process (in its logarithmic form) spends a certain random time below level $\alpha$ after time $L_{\alpha}$. This setup is quite reasonable since, as in other intensity-based models, the market observers are unaware of the firm’s state after a certain time (represented by $L_{\alpha}$ in our model) and receive a default as a sudden shock.
Remark 3.1.
Since the dynamics (3.2) is simple enough, we can compute the distribution of default time $\xi$ explicitly. From (3.4) and (3.10), we see that $\nu(J)=J$ since $Z$ does not hit the region $[\alpha^{*},\infty)$: the occupation time of $(-\infty,\alpha^{*})$ after $L_{\alpha}$ until $\xi$ is equal to $J$. Using the fact that $J$ is distributed exponentially with rate 1 and independent of $L_{\alpha}$ , we have, for $T\in\mathbb{R}_{+}$
$\mathbb{P}_{Y_{0}}(\xi\leq T)$ $=\mathbb{P}_{Y_{0}}(L_{\alpha}\leq T-J)=\int_{0}^{T}\mathbb{P}_{Y_{0}}(L_{\alpha}\leq T-x)e^{-x}{\rm d}x$
$=\int_{0}^{T}\mathbb{P}_{Y_{0}}(0<L_{\alpha}\leq T-x)e^{-x}{\rm d}x+\int_{0}^{T}\mathbb{P}_{Y_{0}}(L_{\alpha}=0)e^{-x}{\rm d}x.$
From (3.2) and (3.3), $L_{\alpha}=\sup\{t:\frac{1}{\sigma}\ln(Y_{t})=\alpha^{*}\}$. The last exit time distribution is known: Proposition 4 in Salminen (1984) and Proposition 3.1 in Egami and Kevkhishvili (2020) provide
$\mathbb{P}_{y}(L_{\alpha}\in{\rm d}u)=\frac{p(u;y,\alpha)}{G_{0}(\alpha,\alpha)}{\rm d}u\quad\text{and}\quad\mathbb{P}_{y}(L_{\alpha}=0)=1-\frac{G_{0}(y,\alpha)}{G_{0}(\alpha,\alpha)}$
where $G_{0}(\alpha,\alpha)$ is available in (2.21). It can be confirmed that $\mathbb{P}_{Y_{0}}(L_{\alpha}=0)=0$ when $Y_{0}\geq\alpha$ by (2.21) and Assumption 1. Since the transition density $p$ for the process $\frac{1}{\sigma}\ln(Y_{t})$, a Brownian motion with drift $M$, is available, we can evaluate the integral in the above equation explicitly
$\mathbb{P}_{Y_{0}}(\xi\leq T)$ $=\int_{0}^{T}\left(\int_{0}^{T-x}\frac{p(u;\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G_{0}(\alpha^{*},\alpha^{*})}{\rm d}u\right)e^{-x}{\rm d}x+\int_{0}^{T}\left(1-\frac{G_{0}(\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G_{0}(\alpha^{*},\alpha^{*})}\right)e^{-x}{\rm d}x$ (3.11)
@@
$\hskip 28.45274pt+\left(1-\frac{G_{0}(\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G_{0}(\alpha^{*},\alpha^{*})}\right)(1-e^{-T}),$

### v4

@@ 3.2. Post-last passage time process
See Appendix A.4. ∎
4. Model Implementation
Now we apply the result in Section 3.2 to our asset-to-debt ratio process $Y$ and its post-last passage time process $Z$. Using (3.6) and the scale function $s(y)=\frac{1}{2M}(1-e^{-2M\cdot y})$ of the Brownian motion with drift $M$, we see that the process $Z_{t}$ follows the dynamics (4.1) right below; that is, the post-$\lambda_{\alpha^{*}}$ dynamics of the normalized asset-to-debt ratio process:
@@ 4. Model Implementation
This is the process we shall work with for explicitly calculating LGD distribution. The starting point of $Z$ is $\alpha^{*}$. Note that since $\lim_{x\uparrow 0}\coth(x)=$-$\infty$, when the process $Z$ approaches $\alpha^{*}$ from below, the drift approaches negative infinity so that the process shall never reach the point $\alpha^{*}$ from the region $(-\infty,\alpha^{*})$. Mathematically, as proved in Proposition 3.3, $\alpha^{*}$ is an entrance boundary for $Z$ and financially, the firm’s asset-to-debt ratio process $Y$ shall never recover to the threshold level $\alpha$.
@@ 4.1. Loss-given-default distribution
The next task is to model the default time $\xi$. We define a random time that is suitable for default time. As is seen in the intensity-based default modeling, we need an exponential random variable $J$ with rate $\eta=1$, independent of everything else. We let $\Gamma:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ be a nonnegative piecewise continuous function and define the moment inverse of integral functional by
$\nu(J):=\min\left(s:\int_{0}^{s}\Gamma\left(Z_{u}\right){\rm d}u=J\right).$ (4.2)
@@
We assume that $\Gamma$ satisfies
$\int_{0}^{\infty}\Gamma\left(Z_{s}\right){\rm d}s=\infty\quad\mathbb{P}-a.s.$ (4.3)
@@
$\xi(\omega):=\lambda_{\alpha}(\omega)+\tau\circ\theta(\lambda_{\alpha}(\omega))$ (4.4)
where $\theta(\cdot)$ denotes the shift operator. Note that the construction of the random time is similar to the canonical construction of the default time in the intensity-based approach (see Proposition 5.26 in Capiński and Zastawniak (2017)). It follows from (4.2) that $\mathbb{P}\left(\tau=0\mid Z_{0}=\alpha^{*}\right)=0$ and from (4.3) that $\mathbb{P}\left(\tau<+\infty\mid Z_{0}=\alpha^{*}\right)=1$. We also see that $\mathbb{P}\left(\tau>s\mid Z_{0}=\alpha^{*}\right)>0$ for every $s\geq 0$. Moreover, we have $\mathbb{P}(\lambda_{\alpha}<\xi<+\infty)=1$ with (4.4). We have constructed the time $\tau$ using the process which never returns to $\alpha^{*}$, this feature being represented in (4.1).
@@
Due to $Y_{\xi}=Y_{\tau}\circ\theta(\lambda_{\alpha})$, it suffices to compute the distribution of $Z_{\tau}=\frac{1}{\sigma}\ln(Y_{\tau}\circ\theta(\lambda_{\alpha}))$. Thus, we are able to find the LGD distribution by focusing on $Z_{\tau}$ only. This simplifies the analysis as we do not need to consider the shift operator $\theta$ in the definition (4.4). Our model is summarized in Table 1: we have made only Assumptions 1 and 2.
@@
Our approach $\xi=\lambda_{\alpha}+\tau\circ\theta({\lambda_{\alpha}})$ $Y_{\xi}\overset{d}{\sim}e^{\sigma Z_{\tau}}$
@@
$U(z):=\mathbb{E}_{z}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right],\quad z<\alpha^{*},\quad\gamma>0$ (4.6)
@@
$\frac{1}{2}U^{\prime\prime}(z)+M\coth(M(z-\alpha^{*}))U^{\prime}(z)-(\Gamma(z)+\gamma)U(z)=-\Gamma(z)\mbox{1}\hskip-2.5pt\mbox{l}_{[-\infty,Q]}(z),\quad z<\alpha^{*}.$ (4.7)
Here $\mathbb{E}_{z}[\cdot]$ denotes the conditional expectation $\mathbb{E}[\cdot\mid Z_{0}=z]$. We will provide an explicit form of $U(z)$ in Proposition 4.2. Using the solution $U(z)$, we obtain $\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]=\lim_{z\uparrow\alpha^{*}}U(z)$ by the continuity of $U(z)$. Finally, $\lim_{\gamma\downarrow 0}\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]$ produces the distribution of $K^{B}(\xi)$ via (4.5).
To obtain a concrete result, we specify the function $\Gamma$ in (4.2) as
$\Gamma\left(Z_{t}\right):=\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{t}<\alpha^{*}\}}$ (4.8)
so that the integral in (4.2) represents the occupation time of the process $Z$ under the level $\alpha^{*}$. It follows that we identify the firm’s default once the asset-to-debt ratio process (in its logarithmic form) spends a certain random time below level $\alpha$ after time $\lambda_{\alpha}$. This specification is quite reasonable and natural since deterioration of credit quality is measured by the duration under water. As in other intensity-based models, the market observers are unaware of the firm’s state after a certain time (represented by $\lambda_{\alpha}$ in our model) and receive a default as a sudden shock.
Since the dynamics (2.5) is simple enough, we can compute the distribution of default time $\xi$ explicitly. From (4.2) and (4.8), we see that $\nu(J)=J$ since $Z$ does not hit the region $\alpha^{*},\infty)$: the occupation time of $(-\infty,\alpha^{*})$ after $\lambda_{\alpha}$ until $\xi$ is equal to $J$. This observation leads to the following result:
Proposition 4.1.
Under (4.8), the distribution of the default time $\xi$ is given by
$\mathbb{P}_{Y_{0}}(\xi\leq T)$ $=\int_{0}^{T}\left(\int_{0}^{T-x}\frac{p(u;\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G(\alpha^{*},\alpha^{*})}{\rm d}u\right)e^{-x}{\rm d}x+\int_{0}^{T}\left(1-\frac{G(\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G(\alpha^{*},\alpha^{*})}\right)e^{-x}{\rm d}x$ (4.9)
@@
$\hskip 28.45274pt+\left(1-\frac{G(\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G(\alpha^{*},\alpha^{*})}\right)(1-e^{-T}),$